OpenStudy (anonymous):

Describe the mathematic process of canceling like factors when working with rational expressions. Demonstrate this with an example.

OpenStudy (anonymous):

x² + 8x + 15 (x + 3)(x + 5) ----------- = -------------- x² - x - 12 (x + 3)(x - 4) By factorising both numerator and denominator we find that (x + 3) is a common factor, so we cancel it and are left with x + 5 ----- x - 4

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

can you explain the process in words for rational expressions>

OpenStudy (anonymous):

I got it! :) thanks though

OpenStudy (anonymous):

do i need to explain anything now???

OpenStudy (anonymous):

for this example you should know how to find factors of a quadratic expression...

OpenStudy (anonymous):

to help you, i searched and found the following site... it gives steps how to factorise a quadratic polynomial.... http://www.teacherschoice.com.au/sample_help_1_alg.htm

OpenStudy (anonymous):

thank you, no I understand. thanks you

OpenStudy (anonymous):

u cud also have case like... 2x + 10 ------- x² - 25 here numerator can be rearranged to 2(x + 5) denominator can be rearranged as (x)² - (5)² = (x + 5) (x - 5) so 2x + 10 2(x + 5) ------- = ------------ x² - 25 (x + 5)(x - 5) so cancelling out the common factor (x + 5), we get 2 ------ x - 5